2. An infinitely long wire that has a radius of 3 cm contains a uniform charge density of 15 μC/m2. Determine the total charge present in a 3.0-m long section of the copper wire and the magnitude of the electric field at a distance of 12 m from the wire.



1
Expert's answer
2022-02-15T09:58:14-0500

"E=\frac q{2\pi\varepsilon r}=3.3\cdot 10^{-10}~\frac NC,"

"q=\sigma l=30~\mu C."